If the difference in 25 deg.C and 5 deg.C is 20 deg.C (∆T = 20 deg.C), what is the difference if these temperature are converted to Kelvin, i.e. what is the ∆T in Kelvin?
Answer #1

As we know that 0 oC = 273 K

25 oC = 25 + 273 = 298 K

5 oC = 5 + 273 = 278 K

So, difference in temperature ∆T = 298 − 278 = 20 K

So, the difference between two temperature values are always equal in oC and K scale. But not those temperatures itself.